OpenStudy (anonymous):

Solve by completing the square: x^2-6x+5=0

OpenStudy (anonymous):

\[x^2 -6x=5 => (x-3)^2 = 14\]

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

When he stopped at: \[x^{2}-6x=5\] He added 9 to both sides of the equals sign to allow him to complete the square on the left hand side. \[x^{2}-6x+9=(x-3)^{2}\]

OpenStudy (anonymous):

This si so confusing, I think I need this step by step. Please help!